# From Local to Global Order: A Theory of Neural Synaptic Balance

### v2

@@ 10. Simulations
Figure 9. Stochastic balancing converges to a unique global balanced state (A-B) Simulations use a deep fully connected autoencoder trained on the MNIST dataset. (C-D) Simulations use deep locally connected network trained on the CFAR10 dataset. (A,C) The weights of the network are initialized randomly and saved. The stochastic balancing algorithm is applied and the resulting balanced weights are denoted by $W_{balanced}$. The stochastic balancing algorithm is applied 1,000 different times. In all repetitions, the weights converge to the same value $W_{balanced}$. (B,D) Stochastic balancing decreases the norm of the weights. (A-D) Shaded areas correspond to one standard deviation around the mean.
@@ 11. Discussion
While the theory of neural synaptic balance is a mathematical theory that stands on its own, it is worth considering some of its possible consequences and applications, at the theoretical, algorithmic, biological, and neuromorphic hardware levels.
@@ 11.2. Algorithms
On the algorithmic side, we do not advocate that balancing algorithms should be used systematically, as SGD may reach balanced states by itself when applied with a regularizer, as we have seen in Section 7. Furthermore, balancing does not change the input-output function of a network and thus alone it cannot improve the performance of a network on the training or test sets. Nevertheless, balancing could be used at least: (1) to initialize a network in a balanced state prior to learning; (2) to check that a network trained to minimize $E+R$ has been properly trained by checking that each neuron is balanced with respect to $R$; (3) to balance a non-properly trained network; or (4) to tweak a network trained with ah $L_{p}$ regularization, towards a state where it is $L_{q}$ regularized (with $q\not=p$) without retraining it. It is also possible that periodically balancing a network during training may facilitate learning in some situations.
@@ 11.3. Biology
The balancing operations are local [5], in the sense that they involve only the pre-and post-synaptic weights of a neuron. Thus, unlike backpropagation, balancing is plausible in a physical neural system, as opposed to a digitally simulated neural system and conceivably balancing could be of interest in neuroscience or in neuromorphic engineering, for instance from the standpoint of learning or memory maintenance. While there is extensive literature in neuroscience on the balance between excitation and inhibition in biological networks [30, 26], there is no evidence in favor or against neuronal synaptic balance in the sense described here. However, there is some evidence for the existence of homeostatic processes that scale the input synaptic weights to regulate the activity of neurons [28, 7, 29]. In addition, there is also evidence that biological neurons can scale their intrinsic excitability (spike threshold) to regulate their activity [14, 11]. It is at least conceivable that these two scaling mechanisms could be at play and work together in some situations. In any case, current technology is quite far from allowing one to measure the strength of all the incoming and outgoing synapses of a biological neuron in an animal brain. Thus, it is impossible to draw any conclusions on the existence of some kind of homeostasis between the incoming and outgoing synapses of biological neurons. Exploratory experiments could potentially be carried either in simpler organisms with a small number of well-defined neurons, such as C. elegans, or in cultured neurons. Finally, simulations could be carried out in neural network models where inhibitory and excitatory neurons are segregated, possibly in uneven proportions–the overall ratio of excitatory to inhibitory neurons in the mammalian cortex is roughly 80% to 20% [21], with biologically-observed connectivity patterns.
@@ 11.4. Neuromorphic Hardware
In physical neural networks (e.g. biological neural networks, neuromorphic chips), as opposed to digitally simulated neural networks, the algorithms for adjusting synaptic weights must be local both in space and time. Yet these networks must exhibit good global properties. Thus, algorithms that are local and lead to global order, such as the balancing algorithms covered in this work, are of particular interest when considering physical, non-simulated, neural systems. For example, while a deep neuron embedded in a physical network may have no way of monitoring the global training error, conceivably it could sense and monitor its degree of balance, since that is an entirely local property. The degree of balance could be used as a local proxy to monitor global learning progress. Furthermore, the physical properties of the underlying hardware could constrain the synaptic weights or the learning algorithms in ways that could require or benefit from some form of neurall synaptic balance.
In particular, optimization strategies for training spiking neural networks with low energy consumption are analyzed in [24] (see also [22]). In particular, they explicitly mention the homogeneous properties of ReLU neurons and investigate how ReLU scaling can influence the number of spikes generated in each layer and the average energy consumption at each layer. Thus there is a direct connection between our work and some of the neuromorphic literature using scaling methods to control spiking rates and energy consumption, and this connection leads to a new questions for future work: how does balancing affects spiking neurons and energy consumption? Under which hardware conditions does balancing leads to optimal energy consumption?
An interesting special case is provided by memristor devices where the output voltage of neurons is calculated using the equation $V^{o}=V^{i}*GRs$ (see Figure 4 in [15]). In this equation $V^{i}$ is the presynaptic spike voltage, $G$ represents the memristor’s conductance (equivalent to synaptic weights), and $Rs$ is the output resistance of the postsynaptic neuron. The power consumption of the memristor is $(V^{i})^{2})G$, so the lower the value of $G$ is, the lower the power consumption of the memristor and minimizing the $L_{2}$ norm is directly connected to minimizing power consumption. Since our balancing algorithm is aimed precisely at minimizing the $L_{2}$ penalty), this establishes a direct connection between neural balance and energy minimization in memristor-based neuromorphic hardware. Moreover, the issue of the limited conductivity range of memristors is mentioned in [15] and in [16] Therefore, reducing the norm of the weights can help mitigate this issue as well.

### v3

@@ 10. Simulations: Corroboration of the Theory
Figure 9. Stochastic balancing converges to a unique global balanced state (A-B) Simulations use a deep fully connected autoencoder trained on the MNIST dataset. (C-D) Simulations use deep locally connected network trained on the CIFAR10 dataset. (A,C) The weights of the network are initialized randomly and saved. The stochastic balancing algorithm is applied and the resulting balanced weights are denoted by $W_{balanced}$. The stochastic balancing algorithm is applied 1,000 different times. In all repetitions, the weights converge to the same value $W_{balanced}$. (B,D) Stochastic balancing decreases the norm of the weights. (A-D) Shaded areas correspond to one standard deviation around the mean.
11. Simulations: Improving Training
The code and experiments for this section are publicly available at https://github.com/antonyalexos/Neural-Balance. In these simulations, we show that full balancing before training, as well as alternating partial balancing with stochastic gradient descent, are effective and practical regularization approaches that often outperform stochastic gradient descent paired with a traditional regularization function. We train all of our models on a server equipped with 8 Nvidia RTX A6000 Ada Generation graphics cards, with 384 GB of total memory, run on CUDA version 12.4. To ensure that our results are reproducible and fair, we repreat the experiments 8 times with 8 different random seeds shared across the different methodologies and take the average. In the experiments we use both $L_{1}$ and $L_{2}$ balancing. To assess both full balance before training and partial balance during training, we combine a variety of neural balancing operations. For the full balancing operation, we apply the balancing operation to the neurons of the model after weight initialization and before training until the ratio of the input to output norms converge to within 0.01 of 1. To assess partial balance during training, we perform partial balancing operations on the neurons of the model at every epoch during training. The partial balancing procedure applies the balancing operation to the neurons in the network only once, usually going from input neurons to output neurons without reaching equilibrium. In all the Tables in this section, optimal results are shown in bold.
Figure 10. Partial neural balancing performed at every epoch on a toy model with 5 neurons, trained on a concentric-circle binary classification dataset. We observe that the norms of the model’s weights gradually converge to 1.
Figure 11. Full neural balancing performed at every epoch on a toy model with 5 neurons, trained on a concentric-circle binary classification dataset. We observe that the norms of the model’s weights converge to 1 almost instantly.
11.1. Toy Experiment on a Circle Toy Dataset
The toy experiments uses a simple 2-dimensional concentric circle clssification class and a simple network containing 2 input neurons, for each coordinate in the 2-dimensional plane, a hidden layer with 5 neurons, and a single output neuron for the binary classification task. The input and hidden layers apply a ReLU activation, and the output neuron uses a logistic activation for classification. The layers are fully connected. The loss is calculated using the binary cross entroyp for classification tasks. We use Adaptive Moment Estimation (Adam) as the optimizer during training, with a learning rate of $0.01$, and we train the model for 1000 epochs. To compare the outcome of full balancing with partial balancing operations, we perform both during training at every epoch. The full balancing procedure applies the balancing operation to the neurons in the network until the ratio of the input and output norms of every neuron converge within .01 of 1. The partial balancing procedure applies the balancing operation to the neurons of the neural network only once every epoch, proceeding from the input to the output neurons.
Figure 11 illustrates the effect of partial balancing applied at each learning epoch and shows that input and output weight norms become equalized for every neuron over time Figure 11 illustrates the effect of full balancing with instantaneous equalization of the input and output norms of all the neurons. While both approaches converge to a balanced state, partial balancing is computationally less intensive and thus provides an effective alternative.
11.2. Full Balancing Before Training Applied to Fully Connected Networks
In these experiments, we study the effect of full balancing prior to training on fully connected networks of various sizes. The networks are trained on the MNIST dataset. The networks are fully connected with an input size of 784, representing the flattened input shape of 28x28, and an output size of 10. The 2 layer model has a single hidden layer with 256 neurons; the 3 layer model has 2 hidden layers with 25 and 128 neurons; and the 5 layer model has 4 hidden layers with 512, 256, 128, and 64 neurons. All neurons have ReLU activation functions. We optimize the cross entropy loss for classification with stochastic gradient descent with a learning rate of 0.001. For experiments using $L_{1}$ and $L_{2}$ regularization, we use a regularization coefficient $\lambda=0.015$ as we found it to be the most effective in this specific experiment via hyperparameter tuning.
In Figure 12 and Table 1, we evaluate the impact of applying the full balancing operation prior to the commencement of training. Compared to standard initialization, full balancing leads to faster convergence and higher overall accuracy when using the same model architecture, hyperparameters, and training methodologies. Notably, the size of the benefit seems to increase with the size of the models. Figure 12 provides a comparison between full balancing and no balancing applied to a Fully Connected Network (FCN) before training. Each square within the grid represents a combination of model size and training methodology, consistently demonstrating that neural balancing enhances the rate of convergence and model accuracy.
Figure 12. The effect of full balancing (FB) before the start of training on various sizes of fully connected networks (FCN) with 2,3, and 5 hidden layers, trained on the MNIST dataset, without regularization, with $L_{1}$ regularization, and with $L_{2}$ regularization. FB at the beginning of learning results in faster convergence and higher test accuracy. The shading in each figure highlights the improvement. Table 1 reports the final text accuracy values.
Type No FB at Start FB at Start
Plain L1 Reg. L2 Reg. Plain L1 Reg. L2 Reg.
2 Layer FCN 90.09% 90.05% 90.062% 91.22% 93.96% 91.18%
3 Layer FCN 89.594% 89.67% 89.70% 90.83% 93.47% 90.79%
5 Layer FCN 89.09% 87.85% 90.3% 91.37% 95.50% 91.59%
Table 1. Test accuracy after training of plain, $L_{1}$-regularized, and $L_{2}$ regularized fully connected networks trained on MNIST, comparing full balancing (FB) before training with no balancing before training. As already observed in Figure 12, full balancing before training results in higher test accuracy.
11.3. Partial Balancing During Training Applied to Fully Connected Networks
Here we conduct the exact same experiments as in the previous section (11.2), but this time we use partial balancing during training in alternation with stochastic gradient descent steps.
From Figure 13 and Table 2, we observe that partial balancing during training results in faster convergence and better test accuracy.
Figure 13. The effect of partial balancing during training on various sizes of fully connected networks (FCN) with 2,3, and 5 hidden layers, trained on the MNIST dataset, without regularization, with $L_{1}$ regularization, and with $L_{2}$ regularization. Partial balancing results in faster convergence and higher test accuracy. Table 2 contains the tabularized data from the plots for easier comparison.
Type Plain L1 PB L2 PB L1 Reg. L2 Reg.
2-FCN 91.22% 94.54% 91.19% 93.96% 91.18%
3-FCN 90.84% 93.94% 90.86% 93.47% 90.79%
5-FCN 91.37% 96.26% 91.63% 95.48% 91.59%
Table 2. Comparison of test accuracy obtained with different training methods and 3 architectures of increasing depth size. As already observed in Figure 13, L1 partial balancing outperforms the other training methodologies on all model sizes.
For increased robustness, we conduct the same tests using the FashionMNIST dataset. We utilize similar fully connected networks of various sizes and implement partial balancing at every epoch. Similar improvements in convergence and performance are observed. Irrespective of model size or training methodology, partial neural balancing markedly enhances the rate of convergence and overall test accuracy. As the models grow bigger, the inclusion of partial neural balancing in training helps models converge faster and perform better when compared to the other techniques.
Figure 14. The effect of partial balancing during training on various sizes of fully connected networks (FCN) with 2,3, and 5 hidden layers, trained on the FashionMNIST dataset. Partial balancing results in faster convergence and higher test accuracy. As the size of the models increases, partial balancing during training helps models converge faster and perform better.
11.4. Full Balancing Before Training Applied to Recurrent Neural Networks
We continue the assessment of synaptic neural balancing by applying it to recurrent neural networks (RNNS). More specifically we assess the effect of full balancing before training. The experiments are carried on the IMDB sentiment analysis dataset. We use a recurrent neural network with an input layer, and output layer, and a fully-connected recurrent hidden layer. Plain text is first passed through an embedding layer, converting it to an embedding vector of length 100. The network uses 256 hidden neurons, which are fully connected among themselves. These recurrent neurons are updated three times synchronously before computing the final output. The final classification output is produced by a single logistic neuron. We use the binary cross entropy loss for training together with stochastic gradient descent with a learning rate of 0.001. For experiments using $L_{1}$ and $L_{2}$ regularization, we use a regularization rate of $\lambda=0.0001$ as we found it to be the most effective via hyperparameter tuning.
The results are shown in Table 3, demonstrating that, in all cases, when full balancing is performed before training, the model has better final accuracy.
Type Plain L1 Regularization L2 Regularization
No FB at Start 88.26 88.23 88.22
FB at Start 88.64% 88.24% 88.57%
Table 3. Test accuracy for a recurrent neural network trained on the IMDB sentiment analysis dataset, comparing plain, $L_{1}$-regularized, and $L_{2}$-regularized models with and without full balancing (FB) at the start of training. In all cases, full balancing before training results in higher test accuracy at the end of training.
11.5. Partial Balancing During Training Applied to Recurrent Neural Networks
Similarly, here we assess the application of partial balancing during training, and demonstrate its efficacy at increasing the speed of convergence and the overall test accuracy, for recurrent neural network architectures. The experiments are also carried out on the IMDB sentiment analysis dataset, using the same recurrent network as in the previous section.
The main results are summarized in Figure 15 where we compare the test accuracy achieved with plain training (no regularization and no balancing), training with partial balancing, and training with $L_{2}$ regularization. We also compare implementing full balancing before the start of training to standard random initialization. Both partial balancing during training and full balancing before training lead to faster convergence and higher accuracy.
Figure 15. Comparison between plain training (no regularization, no partial balancing) training with partial balancing, and training with $L_{2}$ regularization of a recurrent neural network on the IMDB dataset. We also contrast the standard random initialization with full balancing performed before the start of training. Both partial balancing during training, and full balancing before training, result in faster convergence, and higher overall accuracy.
11.6. Neural Balancing with Limited Data Environments
To further examine the reegularizing benefits of synaptic balancing, here we vary the amount of available training data. For these experiments, we use both the MNIST handwritten digit classification dataset and the IMDB sentiment analysis dataset. Unlike the previous experiments, we use a stratified sample of the dataset to simulate a limited-data environment for the models. We use Scikit-Learn to take a stratified split of the overall dataset. For the MNIST dataset, we take 600 samples, representing 1% of the full dataset, with 60 samples from each label, in order to maintain the label balance present in the main dataset. For the IMDB dataset, we take 1250 samples, representing 5% of the full dataset, with 625 samples from the positive and negative labels, to preserve an equal distribution of labels.
For the experiments with the MNIST handwritten digit recognition dataset, we use the same fully connected network architecture used for the experiments that are performed on the whole dataset. We use the cross entropy loss with stochastic gradient descent with a learning rate of 0.001. For experiments using $L_{1}$ and $L_{2}$ regularization, we use a regularization rate $\lambda=0.015$ as we found it to be the most effective via hyperparameter tuning. We test a combination of full balancing and partial balancing before or during training. With full balancing, the balancing operation is performed after initialization and before training until the ratio of the input to output norms of each neuron equalizes within a threshold of .001 of 1. For partial balancing, we balance each neuron once from inputs to output. This partial balancing step is carried out at the end of each training epoch.
For the experiments with the IMDB sentiment analysis dataset, we use the same recurrent neural network architecture that is used for the experiments performed on the whole dataset. We use binary cross entropy loss with stochastic gradient descent with a learning rate of ).001. For experiments using $L_{2}$ regularization, we use a regularization rate $\lambda=0.0001$ as we found it to be the most effective via hyperparameter tuning. We assess the performance of partial balancing by comparing it to $L_{2}$ regularization and plain train (no regularizaation, no balancing). For partial balancing, we balance each neuron once from inputs to output. This partial balancing step is carried out at the end of each training epoch.
Figure 16. Comparison between plain training (no balancing, no regularization), training with partial balancinge, and training with regularization, with and without a full balancing at the start of training, using different fully connected networks trained on only 1% of the MNIST dataset. Table 4 shows the final test accuracy results for each of the training methodologies. Nneural balancing consistently improves both the rate of convergence and the overall accuracy of the model.
Type No FB at Start FB at Start
2 Layer FCN 3 Layer FCN 5 Layer FCN 2 Layer FCN 3 Layer FCN 5 Layer FCN
Plain 84.15% 73.49% 88.9% 91.39% 91.42% 90.86%
L1 PB 91.25% 90.57% 92.87% 93.26% 93.3% 92.92%
L2 PB 87.99% 84.53% 91.06% 91.94% 91.37% 90.59%
L1 Reg. 83.92% 72.03% 11.35% 85.98% 81.33% 19.79%
L2 Reg. 83.35% 75.21% 86.81% 91.16% 90.86% 88.78%
Table 4. Test accuracy of various methodologies trained using 1% of the MNIST dataset to simulate a limited data environment. Consistently with Figure 16, in all cases the use of full balancing at the start of training not only increases the rate of convergence, but also leads to higher accuracy.
In data-scarce environments, neural networks incorporating balancing techniques exhibit accelerated convergence and enhanced accuracy compared to both unmodified models and those employing traditional regularization methods. Both full balancing at the start of learning, and partial balancing during learning lead to notable improvements.
Figure 17. Comparison between plain training (no regularization, no balancing), training with partial balancing, and training with $L_{2}$ regularization using a RNN and only 5% of the IMBD dataset. Partial balancing leads to the best overall performance. Table 5 shows the final test accuracy.
Type Plain L2 Partial Balancing L2 Regularization
87.87% 88.55% 88.34%
Table 5. Test accuracy for a recurrent neural network trained on 5% of the IMDB sentiment analysis dataset, comparing plain training, training with partial balancing, and training with $L_{2}$ regularization. Consistently with the results in Figure 17, $L_{2}$ partial balancing performed at every epoch leads to the highest degree of accuracy.
We continue this assessment by training the same RNN as above using only 5% of the IMBD dataset. The results are provided in Figure 17 and Table 5. Partial balancing leads to faster convergence and higher accuracy, and thus in general balancing can improve generalization in data scarce environments.
11.7. Neural Balance with Non-BiLU Activation functions
In all the previous experiments, we used neurons with ReLU activation functions. Here, we conduct similar experiments with non-BiLU neurons, in particular using sigmoidal neuron ($\tanh$ and logistic). For these experiments, we parallel the full balance experiments in section 11.2, and the partial balancing experiments in section 11.3. For both experiments, we use the same fully connected fedforward architectures with the same sizes, and the same hyperparameter settings as in the aforementioned sections.
Figure 18. Effect of full balancing (FB) before the start of training of fully connected feedforward networks of various depths, with $\tanh$ activation functions, using different training approaches: plain (no regularization, no balancing), training with $L_{1}$ or $L_{2}$ regularization. Regardless of the training method, full balancing at the start results in faster convergence and higher accuracy. The shading illustrates the performance improvement. Table 6 provides the final accuracy values.
Type No FB at Start FB at Start
Plain L1 Reg. L2 Reg. Plain L1 Reg. L2 Reg.
2 Layer FCN 91.69% 87.11% 91.72% 91.97% 89% 91.9%
3 Layer FCN 92.28% 82.75% 92.03% 92.58% 90.23% 92.08%
5 Layer FCN 91.64% 87.61% 90.99% 92.48% 91.28% 91.91%
Table 6. Test accuracy for plain, $L_{1}$-regularized, and $L_{2}$-regularized fully connected feedforward networks with $\tanh$ units trained on the MNIST dataset, comparing the effect of full balancing before training to no full balancing before training. Consistently with the results displayed in in Figure 18, full balancing before training results in faster convergence, as well as higher test accuracy, across all methods.
In figure 18, we observe an improvement in both the rate of convergence and the final accuracy when full balancing is applied at the start of training. This is true across the architectures that were testes and across the training methodologies, with and without regularization. Thus balancing at the start of training appears to be useful even when non-BiLU neurons are used. Note that $\tanh$ units behave like BiLU units, and more precisely like linear units, when the weights are very small, i.e., at the beginning of training.
Figure 19. Comparison of partial $L_{1}$ and $L_{2}$ balancing to $L_{1}$-regularized and $L_{2}$-regularized training for different architectures of $\tanh$ units trained on the MNIST dataset. As the models grow bigger, partial balancing helps the models converge faster and perform better. Table 7 reports the final test accuracy values.
Type Plain L1 PB L2 PB L1 Reg. L2 Reg.
2-FCN 91.69% 93.07% 92.0% 87.11% 91.72%
3-FCN 92.28% 93.66% 92.27% 82.75% 92.03%
5-FCN 91.64% 93.44% 92.36% 87.61% 90.99%
Table 7. Comparison of test accuracy across training methods including plain training (no regularization, no balancing), partial balancing (PB) using $L_{1}$ or $L_{2}$ costs, and $L_{1}$ or $L_{2}$ regularized training using the MNIST dataset for fully connected networks (FCNs) of depth 2, 3, and 5. The networks use $\tanh$ actvation functions As observed in the plots in Figure 19, $L_{1}$ partial balancing outperforms the other training methodologies across all model sizes.
Next we continue with a a comparison of partial balancing with other training methodologies on the MNIST dataset, using $\tanh$ activation functions. We observe through Figure 19 and Table 7 that partial balancing improves convergence and accuracy. In this case, the best results are obtained with $L_{1}$ partial balancing. The use of $L_{1}$ partial balancing significantly improves the rate of convergence and the final test accuracy of the models, irrespective of their size. Again this demonstrates the effectiveness of balancing even in models that use non-BiLU activation functions.
@@ 12. Discussion
While the theory of synaptic neural balance is a mathematical theory that stands on its own, it is worth considering some of its possible consequences and applications, at the theoretical, algorithmic, biological, and neuromorphic hardware levels.
@@ 12.2. Algorithms
The theory and the simulations in Section 11 show that synaptic balance has a number of different applications. It can be used: (1) To initialize the weights of a network at the start of training to better condition the learning; (2) To check that a network trained to minimize $E+R$ has been properly trained by checking that each neuron is balanced with respect to $R$; (3) To balance a non-properly trained network at the end of training; (4) To better regularize a network by alternating SGD steps applied to $E$, or even $E+R$, with partial or full balancing steps. This may improve the convergence or final performance of learning; (5) this of course can be applied to all situations where regularization is beneficial, including the case of smaller training sets; and (6) To tweak a network trained with ah $L_{p}$ regularization, towards a state where it is $L_{q}$ regularized (with $q\not=p$) without retraining it. Thus, in short, balancing is a little bit like dropout: it ought to be added to the arsenal of tools available to improve neural network training.
@@ 12.3. Biology
The balancing operations are local [7], in the sense that they involve only the pre-and post-synaptic weights of a neuron. Thus, unlike backpropagation, balancing is plausible in a physical neural system, as opposed to a digitally simulated neural system. Thus synaptic balancing could be of interest in neuroscience or in neuromorphic engineering, for instance from the standpoint of learning or memory maintenance. While there is extensive literature in neuroscience on the balance between excitation and inhibition in biological networks [35, 31], there is little evidence in favor or against neuronal synaptic balance in the sense described here. However, there is some evidence for the existence of homeostatic processes that scale the input synaptic weights to regulate the activity of neurons [33, 9, 34]. In addition, there is also evidence that biological neurons can scale their intrinsic excitability (spike threshold) to regulate their activity [16, 13]. It is at least conceivable that these two scaling mechanisms could be at play and work together in some situations (see also [30]). In any case, current technology is quite far from allowing one to measure the strength of all the incoming and outgoing synapses of a biological neuron in an animal brain. Thus, it is difficult to draw any definitive conclusions on the existence of some kind of homeostasis between the incoming and outgoing synapses of biological neurons. Exploratory experiments could potentially be carried either in simpler organisms with a small number of well-defined neurons, such as C. elegans, or in cultured neurons. In addition, simulations could be carried out in more detailed compartmental neural network models, possibly ones where inhibitory and excitatory neurons are segregated in uneven proportions–the overall ratio of excitatory to inhibitory neurons in the mammalian cortex is roughly 80% to 20% [25], with biologically-observed connectivity patterns. Finally, as a most speculative and for now untestable hypothesis, one may conjecture that biological neurons may undergo balancing phases, and that perhaps those phases occur during the night, with a close association between synaptic balancing and dreaming.
@@ 12.4. Neuromorphic Hardware
In physical neural networks (e.g. biological neural networks, neuromorphic chips), as opposed to digitally simulated neural networks, the algorithms for adjusting synaptic weights must be local both in space and time. Yet these networks must exhibit good global properties. Thus, algorithms that are local and lead to global order, such as synaptic balancing, are of particular interest when considering physical, non-simulated, neural systems. For example, while a deep neuron embedded in a physical network may have no way of monitoring the global training error, conceivably it could sense and monitor its degree of balance, since that is an entirely local property. The degree of balance could be used as a local proxy to monitor global learning progress. Furthermore, the physical properties of the underlying hardware could constrain the synaptic weights or the learning algorithms in ways that could require or benefit from some form of synaptic balance.
Optimization strategies for training neuromorphic spiking neural networks with low energy consumption are analyzed in [29] (see also [26]), including how the homogeneous properties of ReLU neurons can influence the number of spikes generated at each layer and the average energy consumption at each layer. Thus there is a direct connection between synaptic balance and some of the neuromorphic literature using scaling methods to control spiking rates and energy consumption. In particular, balancing may help minimize energy consumption. Similar considerations apply to memristor-based neuromorphic hardware [17, 18].